How many ways could you hold 5 cards in your hand if 2 of the cards must be aces and the other 3 must be alike?
I have attempted and I get 4P2*(4P3*12) = 3456.
4P2 for the aces and 4P3*12 for the other three cards the 12 because there are twelve other sets of 4 alike cards to choose from. I am using permutations because order matters since it is asking about ways to hold them in your hand??? Is this right??

Answer by stanbon(75887)

How many ways could you hold 5 cards in your hand if 2 of the cards must be aces and the other 3 must be alike?
--------
pick the 2 aces: 4C2 = 6 ways
pick the 3 "alike" cards: 12*4C3 = 12*4 = 48 ways
----------------
Total number of hands: 6*48 = 288
---------------
Total number of permutations of those hands = 5!*288 = 34,560
=============
Cheers,
Stan H.